# Lightning Fast Caching-based Parallel Denoising Prediction for Accelerating Talking Head Generation

### v1

@@ Introduction
Audio-driven talking head generation has received significant attention due to its wide applications in virtual avatars, digital content creation, and real-time communication (Prajwal et al. 2020; Zhang et al. 2023; Wang et al. 2022; Tian et al. 2024). Recent advancements in generative diffusion models (Tian et al. 2024; Xu et al. 2024a; Chen et al. 2025; Zheng et al. 2024) have greatly enhanced the realism and quality of talking head videos, surpassing earlier GAN-based techniques (Prajwal et al. 2020; Wang et al. 2021; Zhang et al. 2023; Wang et al. 2022). However, the computational cost associated with diffusion models—primarily due to iterative denoising steps and large parameter sizes—results in slow inference and hinders real-time deployment (Li et al. 2023). To accelerate diffusion models, previous methods reduce the number of denoising steps through advanced samplers (Song, Meng, and Ermon 2020; Lu et al. 2022) or obtain faster models via pruning (Fang, Ma, and Wang 2023), but these often require expensive retraining.
@@
Figure 2: Average attention scores for foreground noisy latent tokens in the reference attention module ($U_{32}$) of Hallo, showing their attention correlations to (a) foreground and background noisy latent tokens and (b) foreground and background reference features. The $L_{2}$ distance of the background attention output features in the upsampling layer $U_{32}$ between consecutive timesteps: (c) reference attention, (d) audio attention, and (e) temporal attention. FG: foreground. BG: background.
More recently, DeepCache (Ma, Fang, and Wang 2024) proposes a training-free method that caches and reuses intermediate features to minimize redundant computations, achieving notable acceleration. FasterDiffusion (Li et al. 2023) reuses encoder features to parallelize decoding across multiple timesteps. However, these methods do not consider the unique redundancies in talking head generation, limiting their effectiveness in this task. To this end, we observed several critical redundancies present uniquely in talking head generation and propose a task-specific acceleration framework for maximal efficiency improvement. Fig. 1 illustrates high temporal redundancy in the input feature $f_{U_{31}}$ to the final upsampling layer $U_{32}$, as measured by both $L_{2}$ distance and cosine similarity, particularly at intermediate sampling stages. To exploit this, we propose a caching mechanism termed Lightning-fast Caching-based Parallel noise prediction (LightningCP) that significantly improves inference speed without sacrificing quality. LightningCP caches and reuses $f_{U_{31}}$ across consecutive timesteps to bypass the entire encoder and most of the decoder, reducing model inference to merely a single upsampling layer. We further propose to parallelize the denoising predictions of these consecutive timesteps, significantly improving inference throughput.
@@
Diffusion models have gained popularity in recent years for generating talking head videos (Tian et al. 2024; Xu et al. 2024a; Chen et al. 2025; Zheng et al. 2024) due to their superior quality and expressiveness. EMO (Tian et al. 2024) introduced an end-to-end model by extending Stable Diffusion (Rombach et al. 2022) with temporal modules and 3D convolutions (Guo et al. 2023). It integrates audio signal into noisy latent via cross-attention and employs Reference Net (Xu et al. 2024b) to maintain facial identity consistency. Subsequent works have followed and improved upon this framework. Hallo (Xu et al. 2024a) proposes a hierarchical audio-driven visual synthesis module to precisely align audio input with visual outputs. EchoMimic (Chen et al. 2025) concurrently utilizes audio and landmark conditions, allowing flexible generation driven by either modality. MEMO (Zheng et al. 2024) enhances identity consistency through a memory-guided temporal module and generates nuanced emotional expressions with an emotion-aware audio module. Hallo3 (Cui et al. 2024) concurrently explores diffusion transformer (DiT) (Peebles and Xie 2023) architectures, capable of generating dynamic objects and immersive backgrounds. Nevertheless, the slow inference speed of these models continue to hinder their practical deployment.
@@
One approach focuses on developing more efficient diffusion samplers such as DDIM (Song, Meng, and Ermon 2020) and DPM-Solver (Lu et al. 2022), reducing inference calls by approximating multiple denoising steps simultaneously. Another line of work employs pruning to obtain faster compact models (Fang, Ma, and Wang 2023) or distills student models that require fewer denoising steps (Salimans and Ho 2022; Meng et al. 2023). However, these methods incur expensive model training.
ToMe (Bolya and Hoffman 2023) proposes a training-free method to accelerate diffusion inference by removing less informative tokens in self-attentions. However, token reduction’s speedup on talking head generation is limited because the pipeline has many auxiliary computations other than self-attention. Recent approaches also leverage caching strategies to reuse intermediate features, skipping redundant computations. DeepCache (Ma, Fang, and Wang 2024) caches high-level features and efficiently updates only low-level features, but it requires inefficient iterative denoising. FasterDiffusion (Li et al. 2023) enables parallel decoding for multiple timesteps by reusing encoder features, yet it involves costly full decoder inference. In contrast, our proposed method enables lightweight parallel inference, achieving superior acceleration without compromising output quality.
@@
The reverse process, parameterized by a neural network, aims to reconstruct the original latent $z_{0}$ from $z_{t}$ by iteratively removing noise. In latent diffusion models, the DDIM sampling strategy is widely adopted for efficient and deterministic inference. The reverse update at each timestep is given by:
@@
As shown in Fig. 3, diffusion-based talking head generation models typically adopt a UNet architecture, consisting of four encoder (downsampling) blocks $D_{0},D_{1},D_{2},D_{3}$, a mid-block $M$, and four decoder (upsampling) blocks $U_{0},U_{1},U_{2},U_{3}$. In each denoising step, the input latent is passed sequentially through the encoder, mid-block, and decoder for denoising prediction. Since this full forward pass is repeated for every diffusion timestep, the process incurs substantial computation cost, especially with large latent diffusion models.
We observe that for most denoising timesteps in talking head generation, all downsampling operations in the encoder and the majority of the upsampling layers in the decoder are redundant. By caching and reusing high-level decoder features across timesteps, it is feasible to bypass the encoder entirely and all but the final upsampling layer of the decoder. As illustrated in Fig. 1, the input feature $f_{U_{31}}$ to the final layer $U_{32}$ of the last upsampling block exhibits substantial temporal similarity across consecutive denoising steps. Specifically, Fig. 1(a) depicts the feature $L_{2}$ distance between adjacent timesteps, while Fig. 1(b) presents the feature cosine similarity matrix computed across all timesteps. Both metrics confirm that $f_{U_{31}}$ maintains high temporal similarity, especially in the middle timesteps. This insight motivates our approach of caching $f_{U_{31}}$ at selected key timesteps and reusing it in subsequent non-key timesteps, significantly reducing computation cost without compromising the fidelity of the generated output.
@@
Unfortunately, reusing the same input latent $z_{t-1}$ across multiple timesteps t2,,tN{t-2,...,t-N} can lead to significant error accumulation, ultimately degrading video quality. As illustrated in Fig. 4(a), the feature similarity between consecutive input latents decreases substantially over time when using samplers like DDIM (Song, Meng, and Ermon 2020). This compounding discrepancy highlights the risk of not updating the input latents during parallel denoising prediction.
@@
This two-step refinement allows us to recover higher-quality output latents by using more accurate model inputs. In practice, we apply Input Latents Estimation after a designated threshold timestep $t_{\text{thresh}}$. In the early denoising phase ($t<t_{\text{thresh}}$), the input latent $z_{t}$ exhibits high similarity across steps, so parallel prediction without correction suffices. Beyond this threshold, the quality benefit of estimating $\hat{z}_{t-i}$ becomes significant. Our experiments confirm that Input Latents Estimation greatly improves video fidelity, particularly for high-speed samplers such as DDIM (Song, Meng, and Ermon 2020), while incurring negligible extra computation.
@@
Once the updated foreground attention output $A_{\mathcal{F}}$ is obtained, we reconstruct the complete attention feature map by merging it with the cached background output $A_{\mathcal{B}}^{(t^{*})}$ from the most recent key timestep $t^{*}$. The final combined feature is given by:
@@
Method Variants HDTF MEAD FLOPS ($\times 10^{12}$/clip) Latency (s/clip) Speedup
@@
Table 2: Ablation study results on HDTF and MEAD. LCP: Lightning-fast Caching-based Parallel noise prediction. DFA: Decoupled Foreground Attention. rm: reference feature removal.
Method Variants HDTF MEAD FLOPS ($\times 10^{12}$/clip) Latency (s/clip) Speedup
@@
We compare our method with two training-free diffusion acceleration baselines: DeepCache (Ma, Fang, and Wang 2024) and FasterDiffusion (Li et al. 2023). For fair comparison, we use identical key timesteps across all acceleration methods for each model, as timestep selection substantially affects efficiency and output quality. All experiments are run using A800 GPUs. We use a single GPU to test latency of the base model and DeepCache, while using multi-GPUs to test latency of FasterDiffusion and our method.
@@
Table 1 presents results on HDTF (Zhang et al. 2021) and MEAD (Wang et al. 2020) datasets, comparing our method with base models and acceleration baselines (Ma, Fang, and Wang 2024; Li et al. 2023) across multiple diffusion-based talking head models (Xu et al. 2024a; Zheng et al. 2024; Chen et al. 2025). Our method achieves the most inference acceleration, with the highest speedup on Hallo (3.15×) (Xu et al. 2024a) and EchoMimic (3.10×) (Chen et al. 2025). While MEMO (Zheng et al. 2024) and EchoMimic Acc. (Chen et al. 2025) have fewer diffusion steps—limiting the number of non-key timesteps—our method is still the fastest on every model. Furthermore, our approach achieves the lowest FLOPs in all settings, demonstrating superior computational efficiency via targeted optimizations: caching temporally stable decoder features and restricting high-cost attention operations to the dynamic foreground.
@@ Conclusion
In this paper, we introduced Lightning-fast Caching-based Parallel noise prediction (LightningCP), a training-free acceleration framework for diffusion-based talking head generation. By caching and reusing high-level decoder features, our approach enables efficient parallel inference across multiple denoising timesteps. Additionally, Decoupled Foreground Attention (DFA) further enhances inference efficiency by restricting attention to the foreground region. We also employ auxiliary techniques to improve generation quality via input latents estimation and introduce extra speedup via reference feature removal. Extensive experiment validates the superior efficiency and quality of our method, which provides a practical and plug-and-play solution for accelerating diffusion-based talking head generation.

### v2

@@ Introduction
Talking head generation has received significant attention due to its wide applications in virtual avatars, digital content creation, and real-time communication (Prajwal et al. 2020; Zhang et al. 2023; Wang et al. 2022; Tian et al. 2024). Recent advancements in generative diffusion models (Rombach et al. 2022; Tu et al. 2024) have greatly enhanced the realism and quality of talking head videos (Tian et al. 2024; Xu et al. 2024a; Chen et al. 2025; Zheng et al. 2024), surpassing earlier GAN-based techniques (Wang et al. 2021; Zhang et al. 2023; Wang et al. 2022). However, the computational cost associated with diffusion models—primarily due to iterative denoising steps and large parameter sizes—results in slow inference and hinders real-time deployment (Sun et al. 2025, 2024a). To accelerate diffusion models, previous methods employ model pruning (Fang et al. 2023) and distillation (Salimans and Ho 2022; Meng et al. 2023) techniques, but these often require expensive training.
@@
Figure 2: Average attention scores for foreground noisy latent tokens in the reference attention module ($U_{32}$) of Hallo, showing attention correlations to (a) FG and BG noisy latent features and (b) FG and BG reference features. The $L_{2}$ distance of the BG attention output features in the upsampling layer $U_{32}$ between consecutive timesteps: (c) reference attention, (d) audio attention, and (e) temporal attention. FG: foreground. BG: background. std: standard deviation.
More recently, DeepCache (Ma et al. 2024) proposes a training-free method that caches and reuses intermediate features to minimize redundant computations, achieving notable acceleration. FasterDiffusion (Li et al. 2023) reuses encoder features to parallelize decoding across multiple timesteps. However, these methods do not consider the unique redundancies in talking head generation, limiting their effectiveness in this task. To this end, we observed several critical redundancies present uniquely in talking head generation and propose a task-specific acceleration framework for maximal efficiency improvement. Fig. 1 illustrates high temporal redundancy in the input feature $f_{U_{31}}$ to the final upsampling layer $U_{32}$, as measured by both $L_{2}$ distance and cosine similarity, particularly at intermediate sampling stages. To exploit this, we propose a caching mechanism termed Lightning-fast Caching-based Parallel denoising prediction (LightningCP) that significantly improves inference speed without sacrificing quality. LightningCP caches and reuses $f_{U_{31}}$ across consecutive timesteps to bypass the entire encoder and most of the decoder, reducing model inference to merely a single upsampling layer. We further propose to parallelize the denoising predictions of these consecutive timesteps, significantly improving inference throughput.
@@
Diffusion-based talking head generation can be classified into two-stage and end-to-end methods. Two-stage methods (Ma et al. 2023; Liu et al. 2024; Li et al. 2025) predict a sequence of intermediate motion representations (e.g. 3DMM, implicit keypoints) from audio, and then generate talking videos from the predicted motion sequence using pretrained GAN-based generators. Although these methods can achieve real-time inference in some cases (Li et al. 2025), their generation quality is limited by the granularity of the motion representation and the rendering quality of the generator. In contrast, end-to-end methods (Tian et al. 2024; Xu et al. 2024a; Chen et al. 2025; Zheng et al. 2024) leverage the powerful generative priors of latent diffusion model (Rombach et al. 2022) and extend it with 3D convolutions, audio attentions, and temporal modules to generate talking head videos. While these methods can achieve higher generation quality, their slow inference speed hinders practical deployment. To bridge this gap, we propose a training-free approach for end-to-end models that yields substantial speedups while preserving quality, taking a significant step forward in real-world application.
@@
Diffusion models can be accelerated via efficient samplers (Song et al. 2020; Lu et al. 2022), model pruning (Fang et al. 2023), and distillation (Salimans and Ho 2022; Meng et al. 2023). However, efficient samplers do not exploit feature redundancy within diffusion models, while pruning and distillation incur expensive model training.
To address these limitations, training-free methods have gained popularity in recent years. ToMe (Bolya and Hoffman 2023) accelerates self-attention by removing less informative tokens, but cannot accelerate talking head generation significantly due to the more complex pipeline. DeepCache (Ma et al. 2024) caches high-level features and efficiently updates only low-level features, but requires inefficient iterative denoising. FasterDiffusion (Li et al. 2023) enables parallel decoding for multiple timesteps by reusing encoder features, but involves costly full decoder inference. In contrast, our cache reduces inference to only a single decoder layer and enables parallel denoising of multiple non-key timesteps with a novel input latents estimation technique.
Foreground-Background Decoupling
The decoupling of foreground (FG) and background (BG) has been used in various downstream tasks of diffusion models. FG-BG segmentation masks are used to preserve subject identity in consistent generation (Wang et al. 2025) and multi-prompt generation (Cai et al. 2025), as well as to preserve background in image and video editing (Zhu et al. 2025; Shuai et al. 2024; Sun et al. 2024b). In talking head generation, our method leverages the static nature of talking head background and employs masks to accelerate model inference for the first time.
@@
The reverse process, parameterized by a neural network, aims to reconstruct the original latent $z_{0}$ from $z_{t}$ by iteratively removing noise. The DDIM sampling strategy is widely adopted for efficient and deterministic inference, with the reverse update at each timestep as:
@@
As shown in Fig. 3, diffusion-based talking head generation models typically adopt a UNet architecture, with four encoder (downsampling) blocks $D_{0},D_{1},D_{2},D_{3}$, a mid-block $M$, and four decoder (upsampling) blocks $U_{0},U_{1},U_{2},U_{3}$. In each denoising step, the input latent is passed sequentially through the encoder, mid-block, and decoder for denoising prediction. Since this full forward pass is repeated for every diffusion timestep, the process incurs substantial computation cost, especially with large latent diffusion models.
We observe that for most denoising timesteps in talking head generation, all downsampling operations in the encoder and the majority of the upsampling layers in the decoder are redundant. By caching and reusing high-level decoder features across timesteps, it is feasible to bypass the encoder entirely and all but the final upsampling layer of the decoder. As illustrated in Fig. 1, the input feature $f_{U_{31}}$ to the final layer $U_{32}$ of the last upsampling block exhibits substantial temporal similarity across consecutive denoising steps. Specifically, Fig. 1(a) depicts the feature $L_{2}$ distance between adjacent timesteps, while Fig. 1(b) presents the feature cosine similarity matrix computed across all timesteps. Both metrics confirm that $f_{U_{31}}$ maintains high temporal similarity, especially in the middle timesteps. This insight motivates us to cache $f_{U_{31}}$ at selected key timesteps and reuse it in subsequent non-key timesteps, significantly reducing computation cost without compromising the fidelity of the generated output.
@@
Unfortunately, reusing the same input latent $z_{t-1}$ across multiple timesteps t2,,tN{t-2,...,t-N} can lead to significant error accumulation, ultimately degrading video quality. As illustrated in Fig. 4(a), the feature similarity between consecutive input latents decreases substantially over time when using samplers like DDIM (Song et al. 2020). This compounding discrepancy highlights the risk of not updating the input latents during parallel denoising prediction.
@@
This two-step refinement allows us to recover higher-quality output latents by using more accurate model inputs. In practice, we apply Input Latents Estimation after a designated threshold timestep $t_{\text{thresh}}$. In the early denoising phase ($t<t_{\text{thresh}}$), the input latent $z_{t}$ exhibits high similarity across steps, so parallel prediction without correction suffices. Beyond this threshold, the quality benefit of estimating $\hat{z}_{t-i}$ becomes significant. Our experiments confirm that Input Latents Estimation greatly improves video fidelity, particularly for high-speed samplers such as DDIM (Song et al. 2020), while incurring negligible extra computation.
@@
Once the updated foreground attention output $A_{\mathcal{F}}$ is obtained, we reconstruct the complete attention feature map by merging it with the cached background output $A_{\mathcal{B}}^{(t^{*})}\in\mathbb{R}^{L_{b}\times d}$ from the most recent key timestep $t^{*}$, where $L_{b}=L-L_{f}$ denotes the number of background tokens. The final combined feature is given by:
@@
Method Variants HDTF MEAD FLOPS $\downarrow$ ($\times 10^{12}$/clip) Latency $\downarrow$ (s/clip) Speedup $\uparrow$
@@
Method Variants HDTF MEAD FLOPS $\downarrow$ ($\times 10^{12}$/clip) Latency $\downarrow$ (s/clip) Speedup $\uparrow$
@@
Table 2: Ablation study results on HDTF and MEAD. LCP: Lightning-fast Caching-based Parallel denoising prediction. DFA: Decoupled Foreground Attention. rm: reference feature removal.
@@
We compare our method with two training-free diffusion acceleration baselines: DeepCache (Ma et al. 2024) and FasterDiffusion (Li et al. 2023). For fair comparison, we use identical key timesteps across all acceleration methods for each model, as timestep selection substantially affects efficiency and output quality. All experiments are run using A800 GPUs. We use a single GPU to test latency of the base model and DeepCache, while using multi-GPUs to test latency of FasterDiffusion and our method.
@@
Table 1 presents results on HDTF (Zhang et al. 2021) and MEAD (Wang et al. 2020) datasets, comparing our method with base models and acceleration baselines (Ma et al. 2024; Li et al. 2023) across multiple diffusion-based talking head models (Xu et al. 2024a; Zheng et al. 2024; Chen et al. 2025). Our method achieves the most inference acceleration, with the highest speedup on Hallo (3.15×) (Xu et al. 2024a) and EchoMimic (3.10×) (Chen et al. 2025). While MEMO (Zheng et al. 2024) and EchoMimic Acc. (Chen et al. 2025) have fewer diffusion steps—limiting the number of non-key timesteps—our method is still the fastest on every model. Furthermore, our approach achieves the lowest FLOPs in all settings, demonstrating superior computational efficiency via targeted optimizations: caching temporally stable decoder features and restricting high-cost attention operations to the dynamic foreground.
@@ Conclusion
In this paper, we introduce Lightning-fast Caching-based Parallel denoising prediction (LightningCP), a training-free acceleration framework for diffusion-based talking head generation. By caching and reusing high-level decoder features, our approach enables efficient parallel inference across multiple denoising timesteps. Additionally, Decoupled Foreground Attention (DFA) further enhances inference efficiency by restricting attention to the foreground region. We also employ auxiliary techniques to improve generation quality via input latents estimation and introduce extra speedup via reference feature removal. Extensive experiment validates the superior efficiency and quality of our method, which provides a practical and plug-and-play solution for accelerating diffusion-based talking head generation.
Ethical Statement
This research is intended to support the development of ethical virtual humans and creative applications in entertainment and communication. However, as with other generative talking head technologies, there is a risk that the proposed acceleration framework could be misused for malicious purposes such as fraud, impersonation, or defamation (Liu et al. 2025). Moreover, the increasing prevalence of AI-generated media may contaminate public datasets and negatively impact the training of downstream AI models (Tu et al. 2025c, a), potentially leading to degraded or unsafe system behavior.
To mitigate these risks, we recommend that any generated talking head videos be clearly watermarked or labeled as AI generated. We are also committed to assist the ongoing community efforts to develop robust evaluations and safety protocols (Tu et al. 2025b), for promoting safe and trustworthy deployment of generative technologies.
Acknowledgements
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).